OpenStudy (mathmath333):

units digit of 1^1*2^2*3^3......100^100 is zero?

OpenStudy (cwrw238):

what is the units digit of 100^100 ?

OpenStudy (mathmath333):

00000000

OpenStudy (mathmath333):

at leat no modolo here

OpenStudy (cwrw238):

well one 0 will do

ganeshie8 (ganeshie8):

modulo kicks in whenever you deal with remainders, so modulo is there here also :)

OpenStudy (mathmath333):

oh but not necessary right

ganeshie8 (ganeshie8):

units digit = 1^1*2^2*3^3......100^100 mod 10 = (1^1*2^2*3^3......100^99)*10 mod 10 = (1^1*2^2*3^3......100^99)*0 mod 10 = 0 mod 10

ganeshie8 (ganeshie8):

depends on how much you love congruences ;)

OpenStudy (mathmath333):

gr8

ganeshie8 (ganeshie8):

a lil more challenging problem would be to find the units digit of SUM : 1^1 + 2^2 + 3^3+ ...... + 100^100

OpenStudy (mathmath333):

its 1

ganeshie8 (ganeshie8):

wow! how ?

OpenStudy (mathmath333):

1^1 just guessed

OpenStudy (mathmath333):

its opposit to mulitly

ganeshie8 (ganeshie8):

lol work it when free, it will be a nice practice for using congruences :)